a lump of steel weighing 30kg at a temp of 427 degree celcius is dropped into 150kg of oil at 27 degree celsius. The specific heat capacity of the steel and the oil are 0.47 KJ/KgK and 2.5 KJ/KgK


1
Expert's answer
2021-07-12T16:22:27-0400

Gives


"m_s=30kg,T=700K,c_s=0.47KJ\/kg-K"

"m_o=150kg,c_o=2.5KJ\/kg-K,T=300K"

"{(ms\u2206T)}_{s}={(ms\u2206T)}_{o}"


"30\times0.47\times(700-T)=150\times(T-300)\times2.5"

"14.1(700-T)=375(T-300)"

"T=315K"

entropy steel


"(∆s)_s=\smallint_1^2\frac{dQ}{T}=\int_1^2\frac{mc_sdT}{T}=mc_sln\frac{T_2}{T_1}"

"(∆s)_s=30\times0.47\times ln\frac{315}{700}=-11.25kJ/K"

Similarly



"(∆s)_o=150\times2.5\times ln\frac{315}{300}=18.29kJ/K"

"\u2206s_{univese}=\u2206s_s+\u2206s_o=-11.25+18.29=7.03KJ\/K"